A Jaguar XK8 convertible has an eight-cylinder engine. At the beginning of its compression stroke, one of the cylinders contains 499 cm3 of air at atmospheric pressure (1.01×105 Pa) and a temperature of 27.0?C. At the end of the stroke, the air has been compressed to a volume of 46.2 cm3 and the gauge pressure has increased to 2.72×10^6 Pa. Calculate the final temperature Tf
Answer #1
Concepts and reason

The main concept used to solve this problem is ideal gas equation, absolute pressure, and gauge pressure.

Initially, write the ideal gas equation. Then, use the ideal gas equation for initial state and final state and then equate using the fact that the gas is in a sealed container. Finally, rearrange the expression for the final temperature. Later, find the final temperature using the expression of the absolute pressure in terms of atmospheric pressure and the gauge pressure, and then calculate final temperature using the given values of the quantities.

Fundamentals

The expression of the ideal gas equation is,

PV=nRTPV = nRT

Here, P is the pressure, V is the volume, n is the number of moles, R is the gas constant, and t is the temperature.

The expression of the absolute pressure is,

Pabs=Patm+Pgauge{P_{{\rm{abs}}}} = {P_{{\rm{atm}}}} + {P_{{\rm{gauge}}}}

Here, Pabs{P_{{\rm{abs}}}} is the absolute pressure, Patm{P_{{\rm{atm}}}} is the atmospheric pressure, and Pgauge{P_{{\rm{gauge}}}} is the gauge pressure.

The expression of the ideal gas equation is,

PV=nRTPV = nRT

The ideal gas equation when the number of moles is equal is,

PVT=nRPVT=constant\begin{array}{c}\\\frac{{PV}}{T} = nR\\\\\frac{{PV}}{T} = {\rm{constant}}\\\end{array}

The expression of the ideal gas equation for initial state and final state of the system when the number of moles is same is,

P1V1T1=P2V2T2\frac{{{P_1}{V_1}}}{{{T_1}}} = \frac{{{P_2}{V_2}}}{{{T_2}}}

Here, P1{P_1} is the initial pressure, V1{V_1}is the initial volume, T1{T_1} is the initial temperature, P2{P_2} is the final pressure, V2{V_2} is the final volume, and T2{T_2} is the final temperature.

Rearrange the expression for the final temperature T2{T_2} at the end of the compression stroke is,

T2=P2V2P1V1T1{T_2} = \frac{{{P_2}{V_2}}}{{{P_1}{V_1}}}{T_1}

The final pressure after the stroke is equal to the gauge pressure and the atmospheric pressure which is given by,

P2=Patm+Pgauge{P_{\rm{2}}} = {P_{{\rm{atm}}}} + {P_{{\rm{gauge}}}}

Substitute 1.01×105Pa1.01 \times {10^5}{\rm{ Pa}} for Patm{P_{{\rm{atm}}}} and 2.72×106Pa2.72 \times {10^6}{\rm{ Pa}} for Pgauge{P_{{\rm{gauge}}}}.

P2=1.01×105Pa+2.72×106Pa=2.82×106Pa\begin{array}{c}\\{P_{\rm{2}}} = 1.01 \times {10^5}{\rm{ Pa}} + 2.72 \times {10^6}{\rm{ Pa}}\\\\ = 2.82 \times {10^6}{\rm{ Pa}}\\\end{array}

The expression of the final temperature after the stroke is,

T2=P2V2P1V1T1{T_2} = \frac{{{P_2}{V_2}}}{{{P_1}{V_1}}}{T_1}

Substitute 1.01×105Pa1.01 \times {10^5}{\rm{ Pa}} for P1{P_1}, 2.82×106Pa2.82 \times {10^6}{\rm{ Pa}} for P2{P_{\rm{2}}}, 499cm3499{\rm{ c}}{{\rm{m}}^3} for V1{V_1}, 46.2cm346.2{\rm{ c}}{{\rm{m}}^3} for V2{V_2}, and 27.0C27.0^\circ {\rm{ C}} for T1{T_1}.

T2=(2.82×106Pa)(1.01×105Pa)(46.2cm3)(499cm3)[(27.0C+273)K]=775.515K\begin{array}{c}\\{T_2} = \frac{{\left( {2.82 \times {{10}^6}{\rm{ Pa}}} \right)}}{{\left( {1.01 \times {{10}^5}{\rm{ Pa}}} \right)}}\frac{{\left( {46.2{\rm{ c}}{{\rm{m}}^3}} \right)}}{{\left( {499{\rm{ c}}{{\rm{m}}^3}} \right)}}\left[ {\left( {27.0^\circ {\rm{ C}} + 273} \right){\rm{ K}}} \right]\\\\ = 775.515{\rm{ K}}\\\end{array}

The temperature in Celsius is,

T2=775.515K273K=502.5C\begin{array}{c}\\{T_2} = 775.515{\rm{ K}} - 273{\rm{ K}}\\\\ = {\rm{502}}{\rm{.5}}^\circ {\rm{ C}}\\\end{array}

Ans:

The final temperature is 502.5C{\rm{502}}{\rm{.5}}^\circ {\rm{ C}}.